# Polyak Minorant Method for Convex Optimization

### v3

@@ 1.5 Prior work
Boris Polyak introduced the Polyak gradient step for minimizing continuously differentiable functionals in [Pol63]. Following the development of subgradient methods for non-differentiable optimization [Sho12, Roc81], Polyak adapted his gradient step to the subgradient case [Pol87]. Since then various extensions of the Polyak subgradient method have been developed, including Polyak step variations for stochastic gradient descent [LVHLLJ21, AXKT23, BZK21, GBGP22, ${}^{+}$23, PO21, OLJL22], Polyak-like steps that do not require knowledge of $f^{\star}$ [HK22, YL22], a Polyak-like step for momentum-accelerated gradient descent [WJZ23, BTd20, GTD22], a Polyak step method for mirror descent [YCL22], a Polyak-like step for convex problems with box constraints [CL12], and a Polyak step for weakly convex functions [DDMP18]. We will see that PMM reduces to the subgradient method with Polyak step sizes when there are no constraints and a subgradient-based affine minorant is used.
@@ 2 Polyak minorant method
where $\mathcal{I}$ is the $\{0,\infty\}$-indicator function. The projection of $x^{k}$ onto $\mathcal{X}^{k}$ also minimizes $t^{k}\hat{F}^{k}(x)+(1/2)\|x-x^{k}\|_{2}^{2}$ for some $t^{k}>0$. In other words, we have $x^{k+1}=\mathop{\bf prox}_{t^{k}\hat{F}^{k}}(x^{k})$, where $\mathop{\bf prox}$ is the proximal operator [${}^{+}$14]. In standard proximal operator methods, $t^{k}$ is specified. In our case, however, $x^{k+1}$ is given as a projection, and we determine $t^{k}$ only after the update is computed.
@@ 2.3 Convergence with sharpness
$\phi(x)-\phi^{\star}\geq\mu\inf_{x^{\star}\in\mathcal{X}^{\star}}\|x-x^{\star}\|_{2},$
@@ 2.4 Cost of an iteration
To illustrate this, we consider a specific example with $n=10^{6}$ and $p=q=50$. Computing the matrices (13) costs $O(10^{10})$ flops, whereas solving the small QP (12) costs $O(10^{6})$ flops, which is negligible in comparison. Carrying out this computation using CVXPY [DB16] and OSQP [${}^{+}$20] on an instance of this problem yields results that are compatible with these rough flop counts. Computing the matrices requires around 0.3 seconds, and solving the small QP requires 0.007 seconds. In comparison, solving the original QP directly takes around 700 seconds. (These numbers are for a laptop with a Ryzen 9 5900HX processor.)

### v4

@@ 1.5 Prior work
Boris Polyak introduced the Polyak gradient step for minimizing continuously differentiable functionals in [Pol63]. Following the development of subgradient methods for non-differentiable optimization [Sho12, Roc81], Polyak adapted his gradient step to the subgradient case [Pol87]. Since then various extensions of the Polyak subgradient method have been developed, including Polyak step variations for stochastic gradient descent [LVHLLJ21, AXKT23, BZK21, GBGP22, LST+23, PO21, OLJL22], Polyak-like steps that do not require knowledge of $f^{\star}$ [HK22, YL22], a Polyak-like step for momentum-accelerated gradient descent [WJZ23, BTd20, GTD22], a Polyak step method for mirror descent [YCL22], a Polyak-like step for convex problems with box constraints [CL12], and a Polyak step for weakly convex functions [DDMP18]. We will see that PMM reduces to the subgradient method with Polyak step sizes when there are no constraints and a subgradient-based affine minorant is used.
@@ 2 Polyak minorant method
where $\mathcal{I}$ is the $\{0,\infty\}$-indicator function. The projection of $x^{k}$ onto $\mathcal{X}^{k}$ also minimizes $t^{k}\hat{F}^{k}(x)+(1/2)\|x-x^{k}\|_{2}^{2}$ for some $t^{k}>0$. In other words, we have $x^{k+1}=\mathop{\bf prox}_{t^{k}\hat{F}^{k}}(x^{k})$, where $\mathop{\bf prox}$ is the proximal operator [PB14]. In standard proximal operator methods, $t^{k}$ is specified. In our case, however, $x^{k+1}$ is given as a projection, and we determine $t^{k}$ only after the update is computed.
@@ 2.3 Convergence with sharpness
$\phi(x)-\phi^{\star}\geq\mu\inf_{x^{\star}\in\mathcal{X}^{\star}}\|x-x^{\star}\|_{2}.$
@@ 2.4 Cost of an iteration
To illustrate this, we consider a specific example with $n=10^{6}$ and $p=q=50$. Computing the matrices (13) costs $O(10^{10})$ flops, whereas solving the small QP (12) costs $O(10^{6})$ flops, which is negligible in comparison. Carrying out this computation using CVXPY [DB16] and OSQP [SBG+20] on an instance of this problem yields results that are compatible with these rough flop counts. Computing the matrices requires around 0.3 seconds, and solving the small QP requires 0.007 seconds. In comparison, solving the original QP directly takes around 700 seconds. (These numbers are for a laptop with a Ryzen 9 5900HX processor.)